Solve: 4/3x -5/6x = 5/6
Answer by edjones(7311)

(4/3)x -(5/6)x = 5/6
multiply 6 times each side: 8x-5x=5
3x=5
divide 3 into each side: x=5/3
Check:
4/3*5/3-5/6*5/3=5/6
20/9-25/18=5/6
40/18-25/18=5/6
15/18=5/6
5/6=5/6
Ed